# RG-Invariant Symmetry Ratio for QCD: A Study of $U(1)_{A}$ and Chiral Symmetry Restoration

### v1

@@ 2.4 Symmetry Channels Involving Flavor Singlet Operators
    In contrast to the scalar and pseudoscalar singlets, the tensor singlet operators $T^{s}_{k}=\bar{q}\gamma_{4}\gamma_{k}q$ and $X^{s}_{k}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ avoid power-divergent mixings and anomalous scale dependence. Their renormalized susceptibilities satisfy
    $\chi_{T}^{R}=(Z_{T}^{s})^{2}\,\chi_{T}^{\text{bare}},\qquad\chi_{X}^{R}=(Z_{X}^{s})^{2}\,\chi_{X}^{\text{bare}},$
    with $Z_{T}^{s}=Z_{X}^{s}$ in chirally symmetric formulations. Consequently, the ratio
    $\kappa_{TX}^{s}=\frac{\chi_{T}^{s}-\chi_{X}^{s}}{\chi_{T}^{s}+\chi_{X}^{s}}$
    is RG‑invariant and provides a theoretically clean probe of $U(1)_{A}$ restoration in the singlet sector. Similarly, the mixed ratios $\kappa(\chi_{T}^{s},\chi_{X}^{ns})$ and $\kappa(\chi_{X}^{s},\chi_{T}^{ns})$ probe $SU(2)_{L}\times SU(2)_{R}$ restoration between singlet and nonsinglet tensor partners.
    The key advantages of tensor singlet channels are:
    • No power divergences: Unlike scalar singlet operators, tensor operators do not mix with the identity operator, avoiding power-divergent subtractions.
    • No anomalous running: The renormalization of tensor operators is purely multiplicative, without the scale-dependent mixing coefficients that plague pseudoscalar singlets.
    • Chiral partner structure: The $T^{s}$ and $X^{s}$ operators are related by $U(1)_{A}$ rotations, while mixed pairs $(T^{s},X^{ns})$ and $(X^{s},T^{ns})$ are related by $SU(2)_{L}\times SU(2)_{R}$ transformations.
    • Practical feasibility: Although still requiring disconnected diagrams, the absence of power divergences and anomalous mixing makes tensor singlet channels theoretically cleaner targets for future high-precision studies.
    Future studies with enhanced computational resources and improved techniques for disconnected diagrams may enable reliable investigation of these channels, offering a more complete view of chiral and axial symmetry restoration across both singlet and nonsinglet flavor sectors.
@@
  • $\kappa_{TX}$ ($\rho$-$b_{1}$) for $U(1)_{A}$ through a different Dirac structure.
These choices provide a clean, theoretically unambiguous, and statistically precise set of observables. The consistent picture emerging from these three channels (presented in sec. 4 and sec. 5) gives strong evidence for the relative restoration scales of $U(1)_{A}$ and $SU(2)_{A}$ symmetries.
@@ 6 Discussion and Outlook
Within a renormalization-group framework, the observed continuum-limit degeneracy of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ implies that $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetry-breaking effects for connected correlators become comparably suppressed near $T_{c}$. This suggests that the corresponding symmetry-breaking operators acquire similar infrared relevance for nonsinglet observables. However, as we discuss in Sec. 6.1, the full effective restoration pattern—including singlet channels—reveals a more intricate structure that refines simple RG scenarios based solely on nonsinglet data.
@@ 6.1 Connections to topological susceptibility and singlet sector
    • $U(1)_{A}$: $\chi_{\pi}\approx\chi_{\delta}$ (Eq. 12), $\chi_{\rho}\approx\chi_{b_{1}}$ (Eq. 14)
@@
  • Stage 2 ($T\sim T_{c}^{s}\gg T_{c}^{ns}$): Full effective restoration including singlet and mixed singlet-nonsinglet channels, requiring $\chi_{t}\to 0$, $\kappa_{TX}^{s}\to 0$ and satisfaction of all relations (eqs. (7)–(15)) including eqs. (25) and (32):
    • $SU(2)_{L}\times SU(2)_{R}$: $\chi_{\sigma}=\chi_{\pi}$, $\chi_{\eta}=\chi_{\delta}$, $\chi_{\omega}=\chi_{b_{1}}$, $\chi_{h_{1}}=\chi_{\rho}$
    • $U(1)_{A}$: $\chi_{\sigma}=\chi_{\eta}$, $\chi_{\omega}=\chi_{h_{1}}$
Figure 4: Schematic illustration of the two-stage hierarchical restoration scenario for light $(u,d)$ quarks. Stage 1 (nonsinglet restoration) occurs around $T_{c}^{ns}\sim 156$ MeV, while Stage 2 (full effective restoration including singlets) requires $\chi_{t}\to 0$ and $\kappa_{TX}^{s}\to 0$ at a much higher temperature $T_{c}^{s}\gg T_{c}^{ns}$. The $\kappa_{AB}$ ratios for nonsinglet channels ($\kappa_{PS}$, $\kappa_{VA}$, $\kappa_{TX}$) probe Stage 1, while tensor singlet ratios ($\kappa_{TX}^{s}$) and direct $\chi_{t}$ measurements probe Stage 2.
@@ 6.2 Hierarchical restoration of chiral and axial symmetries
Notably, in lattice QCD with exact chiral symmetry, $\chi_{t}(T)$ for physical $N_{f}=2+1$ and $N_{f}=2+1+1$ QCD can be much larger than that of $N_{f}=2$ QCD in the chiral limit (see appendix A of ref. Mao:2009sy ). This raises an intriguing possibility: in the $N_{f}=2$ chiral limit, the temperature $T_{c}^{\mathrm{s}}$ for full effective restoration (including singlet channels) could be much closer to $T_{c}^{\mathrm{ns}}\sim T_{c}$ than in the physical $N_{f}=2+1+1$ case. In this hypothetical scenario, $T_{c}^{\mathrm{s}}\gtrsim T_{c}^{\mathrm{ns}}\sim T_{c}$, meaning that the two-stage hierarchy would nearly collapse. Our data do not rule out this possibility, as our simulations are performed at physical masses where the strange and charm quarks explicitly break the symmetry and enhance topological fluctuations.
@@
While $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ cleanly signal nonsinglet restoration near $T_{c}^{ns}$, the singlet sector remains numerically challenging. As discussed in Sec. 2.4, the tensor singlet channels $T^{s}_{k}=\bar{q}\gamma_{4}\gamma_{k}q$ and $X^{s}_{k}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ offer a theoretically cleaner alternative to scalar/pseudoscalar singlets as they avoid power-divergent mixings and possess purely multiplicative renormalization. Measuring the singlet tensor ratio $\kappa_{TX}^{s}\equiv(\chi_{T}^{s}-\chi_{X}^{s})/(\chi_{T}^{s}+\chi_{X}^{s})$ in future studies will therefore provide a direct, renormalization-group invariant probe of full $U(1)_{A}$ effective restoration in the singlet sector. If the hierarchical picture is correct, $\kappa_{TX}^{s}$ should remain non-zero well above $T_{c}^{ns}$ and approach zero only at a higher temperature $T_{c}^{s}\gg T_{c}^{ns}$.
@@ 6.3 Broader symmetry landscape and emergent phases above $T_{c}$
The hierarchical picture presented above aligns with a growing body of evidence for multi-stage transitions above $T_{c}$. Beyond the familiar $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, recent lattice studies reveal emergent approximate symmetries that are absent in the classical QCD Lagrangian. In particular, the approximate chiral-spin $SU(2)_{\mathrm{CS}}$ symmetry Glozman:2014mka observed in $N_{f}=2$ Rohrhofer:2019qwq ; Rohrhofer:2019qal , $N_{f}=2+1+1$ Chiu:2023hnm ; Chiu:2024jyz , and $N_{f}=2+1+1+1$ Chiu:2024bqx lattice QCD suggests the existence of hadron-like states bound predominantly by chromoelectric interactions up to temperatures well above $T_{c}$. The symmetry remains observable even beyond $2T_{c}$, e.g. up to $T\gtrsim 500\;\text{MeV}$, indicating that correlated, hadron-like degrees of freedom survive far into the nominally deconfined regime.
Similarly, the observation of an infrared symmetric phase Alexandru:2019gdm around $T_{\mathrm{IR}}\sim 230$ MeV in $N_{f}=2+1$ lattice QCD Meng:2023nxf , which disappears by $T\gtrsim 300$ MeV, further supports the existence of loosely bound, hadron-like modes in this intermediate temperature window. The temperature range $T_{c}\lesssim T\lesssim T_{\mathrm{IR}}$ is especially intriguing: as $T$ increases above $T_{c}$, hadron-like objects become progressively less bound—yet emergent symmetries such as chiral-spin and the infrared phase indicate that residual chromoelectric binding persists. This suggests a smooth crossover from tightly bound hadrons to a correlated, quasi-hadronic medium, rather than an abrupt dissolution into free quarks and gluons. The gradual loss of binding is consistent with the observed reduction in $U(1)_{A}$-breaking effects in connected channels (as measured by $\kappa_{PS}$ and $\kappa_{TX}$), while topological fluctuations—including center vortices tHooft:1977nqb and instanton-like objects—remain active until higher temperatures.
@@
These emergent symmetry patterns are complemented by topological studies. The persistence of topological effects above the chiral transition is further supported by recent center-vortex analyses in $N_{f}=2+1$ QCD, which observe a pronounced change in vortex percolation and clustering behavior at approximately $\gtrsim 2T_{c}$ Mickley:2024vkm . While the vortex density drops substantially in this region, the survival of a sparse vortex network up to even higher temperatures reinforces the view that topological fluctuations—which are intimately connected to the $U(1)_{A}$ anomaly—remain non-negligible well into the deconfined regime. This microscopic picture provides a plausible mechanism for the continued breaking of $U(1)_{A}$ symmetry at temperatures where the nonsinglet chiral symmetry has been largely restored. Crucially, the smooth nature of the crossover implies that no sharp “complete deconfinement” point exists; instead, the system exhibits a gradual evolution from a hadronic to a quark–gluon plasma phase, with residual correlations and rare, very loosely bound hadron-like states possible even at $T\gg 2T_{c}$.
@@
Taken together, these independent lines of evidence support the two-stage effective restoration scenario described in Sec. 6.1. Stage 1 ($T\sim T_{c}^{ns}$) is characterized by the effective restoration of nonsinglet symmetries, as demonstrated by the vanishing of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ in this work. Stage 2 ($T\sim T_{c}^{s}$) involves the full effective restoration of singlet channels and the suppression of topological fluctuations ($\chi_{t}\to 0$). A consistent comparison between different lattice studies requires continuum extrapolation of each observable, as demonstrated in this work. Since the chiral transition at physical quark masses is a smooth crossover, the value of any symmetry-breaking observable decreases gradually with temperature. The Stage 2 scenario remains to be conclusively established through future continuum-extrapolated measurements of tensor-singlet ratios and topological susceptibility.
@@ 6.4.1 Computational improvements and tensor singlet probes
  • Flavor-singlet tensor channels: Using stochastic all-to-all propagators with color–Dirac dilution, we will compute the singlet tensor ($T^{s}$) and axial-tensor ($X^{s}$) susceptibilities. The tensor singlet ratio
    $\kappa_{TX}^{s}\equiv\frac{\chi_{T}^{s}-\chi_{X}^{s}}{\chi_{T}^{s}+\chi_{X}^{s}},$
    where $T^{s}_{k}=\bar{q}\gamma_{4}\gamma_{k}q$ and $X^{s}_{k}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$, provides a theoretically clean (power-divergence free) probe of $U(1)_{A}$ restoration in the singlet sector. This will enable a direct test of the two-stage $U(1)_{A}$ restoration scenario at $T_{c}^{s}\gg T_{c}^{ns}$.
  • Mixed singlet–nonsinglet channels: The mixed ratios $\kappa(\chi_{T}^{s},\chi_{X}^{ns})$ and $\kappa(\chi_{X}^{s},\chi_{T}^{ns})$ probe $SU(2)_{L}\times SU(2)_{R}$ restoration between singlet and nonsinglet tensor operators, offering additional insight into the chiral symmetry realization pattern across singlet and nonsinglet flavor sectors.

### v2

@@ 2.4 Symmetry Channels Involving Flavor Singlet Operators
    In contrast to the scalar and pseudoscalar singlets, the tensor singlet operators $\omega_{T,k}\equiv T_{k}^{s}=\bar{q}\gamma_{4}\gamma_{k}q$ and $h_{1T,k}\equiv X_{k}^{s}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ avoid both power-divergent mixings and finite mixing with $\chi_{t}$. Their renormalization is purely multiplicative with $Z_{T}^{s}=Z_{X}^{s}$ in chirally symmetric formulations, making the ratio $\kappa_{\omega_{T},h_{1T}}=(\chi_{\omega_{T}}-\chi_{h_{1T}})/(\chi_{\omega_{T}}+\chi_{h_{1T}})$ RG-invariant and theoretically clean. Mixed ratios $\kappa(\chi_{\omega_{T}},\chi_{b_{1}})$ and $\kappa(\chi_{h_{1T}},\chi_{\rho_{T}})$ similarly probe $SU(2)_{L}\times SU(2)_{R}$ restoration between singlet and nonsinglet tensor partners. Future studies with improved techniques for disconnected diagrams may enable reliable investigation of these channels.
@@
  • $\kappa_{TX}$ ($\rho_{T}$-$b_{1}$) for $U(1)_{A}$ through a different Dirac structure.
These choices provide a clean, theoretically unambiguous, and statistically precise set of observables. The consistent picture emerging from these three channels (presented in section 4 and section 5) gives strong evidence for the relative restoration scales of $U(1)_{A}$ and $SU(2)_{L}\times SU(2)_{R}$ symmetries.
@@ 6 Discussion and Outlook
Within a renormalization-group framework, the observed continuum-limit degeneracy of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ implies that $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetry-breaking effects for connected correlators become comparably suppressed near $T_{c}$. This suggests that the corresponding symmetry-breaking operators acquire similar infrared relevance for nonsinglet observables. However, as we discuss in section 6.1, the full effective restoration pattern—including singlet channels—reveals a more intricate structure that refines simple RG scenarios based solely on nonsinglet data.
@@ 6.1 Connections to topological susceptibility and singlet sector
    • $U(1)_{A}$: $\chi_{\pi}\approx\chi_{\delta}$ (Eq. 12), $\chi_{\rho_{T}}\approx\chi_{b_{1}}$ (Eq. 14)
@@
  • Stage 2 ($T\sim T_{c}^{s}\gg T_{c}^{ns}$): Full effective restoration including singlet and mixed singlet-nonsinglet channels, requiring $\chi_{t}\to 0$, $\kappa_{\omega_{T},h_{1T}}\to 0$ and satisfaction of all relations (eqs. (7)–(15)) including eqs. (23) and (30):
    • $SU(2)_{L}\times SU(2)_{R}$: $\chi_{\sigma}=\chi_{\pi}$, $\chi_{\eta}=\chi_{\delta}$, $\chi_{\omega_{T}}=\chi_{b_{1}}$, $\chi_{h_{1T}}=\chi_{\rho_{T}}$
    • $U(1)_{A}$: $\chi_{\sigma}=\chi_{\eta}$, $\chi_{\omega_{T}}=\chi_{h_{1T}}$
Figure 4: Schematic illustration of the two-stage hierarchical restoration scenario for light $(u,d)$ quarks. Stage 1 (nonsinglet restoration) occurs around $T_{c}^{ns}\sim 156$ MeV, while Stage 2 (full effective restoration including singlets) requires $\chi_{t}\to 0$ and $\kappa_{\omega_{T},h_{1T}}\to 0$ at a much higher temperature $T_{c}^{s}\gg T_{c}^{ns}$. The $\kappa_{AB}$ ratios for nonsinglet channels ($\kappa_{PS}$, $\kappa_{VA}$, $\kappa_{TX}$) probe Stage 1, while tensor singlet and mixed singlet-nonsinglet tensor ratios ($\kappa_{\omega_{T},h_{1T}}$, $\kappa(\chi_{\omega_{T}},\chi_{b_{1}})$, $\kappa(\chi_{h_{1T}},\chi_{\rho_{T}})$) and direct $\chi_{t}$ measurements probe Stage 2.
@@ 6.2 Hierarchical restoration of chiral and axial symmetries
Notably, in lattice QCD with exact chiral symmetry, $\chi_{t}(T)$ for physical $N_{f}=2+1$ and $N_{f}=2+1+1$ QCD can be much larger than that of $N_{f}=2$ QCD in the chiral limit (see appendix A of ref. Mao:2009sy ). This raises an intriguing possibility: in the $N_{f}=2$ chiral limit, the two-stage hierarchy might nearly collapse, with $T_{c}^{\mathrm{s}}\gtrsim T_{c}^{\mathrm{ns}}\sim T_{c}$, implying that full effective restoration (including singlet channels) could occur much closer to the chiral transition temperature. Our simulations at physical masses cannot address this scenario directly, as the strange and charm quarks explicitly break the symmetry and enhance topological fluctuations.
@@
While $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ cleanly signal nonsinglet restoration near $T_{c}^{ns}$, the singlet sector remains numerically challenging. As discussed in section 2.4, the tensor singlet operators $\omega_{T,k}\equiv T_{k}^{s}=\bar{q}\gamma_{4}\gamma_{k}q$ and $h_{1T,k}\equiv X_{k}^{s}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ offer a theoretically cleaner alternative to scalar/pseudoscalar singlets. For chirally symmetric fermions, their renormalization constants satisfy $Z_{T}^{s}=Z_{X}^{s}$, and after contact term and vacuum subtraction, their susceptibilities are multiplicatively renormalizable with the same factor $(Z_{T}^{s})^{2}$, ensuring that $\kappa_{TX}^{s}=(\chi_{T}^{s}-\chi_{X}^{s})/(\chi_{T}^{s}+\chi_{X}^{s})$ is RG-invariant. Measuring this ratio in future studies will therefore provide a direct probe of full $U(1)_{A}$ effective restoration in the singlet sector. If the hierarchical picture is correct, $\kappa_{\omega_{T},h_{1T}}$ should remain non-zero well above $T_{c}^{ns}$ and approach zero only at a higher temperature $T_{c}^{s}\gg T_{c}^{ns}$.
@@ 6.3 Broader symmetry landscape and emergent phases above $T_{c}$
The hierarchical picture presented above aligns with growing evidence for multi-stage transitions above $T_{c}$. Beyond the familiar $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, recent lattice studies reveal emergent approximate symmetries absent in the classical QCD Lagrangian. In particular, approximate chiral-spin $SU(2)_{\mathrm{CS}}$ symmetry Glozman:2014mka observed in $N_{f}=2$ Rohrhofer:2019qwq ; Rohrhofer:2019qal , $N_{f}=2+1+1$ Chiu:2023hnm ; Chiu:2024jyz , and $N_{f}=2+1+1+1$ Chiu:2024bqx lattice QCD indicates that hadron-like states bound predominantly by chromoelectric interactions persist well above $T_{c}$.
It is important to distinguish this emergent symmetry from the usual chiral symmetries. For $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$, symmetry breaking decreases monotonically with temperature, and the symmetries themselves become effectively restored as partner degeneracy improves. These symmetries are always present in the Lagrangian — only their breaking weakens.
In contrast, $SU(2)_{\mathrm{CS}}$ chiral-spin symmetry is not present in the QCD Lagrangian. It emerges only in an intermediate window where chromoelectric binding dominates, typically from $T\sim 2T_{c}$ up to $T\sim 4T_{c}$ for light $(u,d)$ quarks. The window boundaries depend on two thresholds: $\epsilon_{\mathrm{CS}}$ for emergence at $T_{\mathrm{CS}}$, and $\epsilon_{\mathrm{CSF}}$ for fading at $T_{\mathrm{CSF}}>T_{\mathrm{CS}}$. Within this window, $SU(2)_{\mathrm{CS}}$ breaking decreases with $T$, allowing hadron-like states bound by chromoelectric interactions to exist. As temperature increases, thermal excitation energy eventually exceeds the chromoelectric binding energy, leading to quark deconfinement. This is manifested when one $SU(2)_{\mathrm{CS}}$ multiplet $(V_{k},T_{k},X_{k})$ merges with the $(P,S)$ multiplet of $U(1)_{A}$ symmetry, leaving only the usual chiral symmetries relevant. At this point, chiral-spin symmetry fades away — not simply becoming more broken, but physically disappearing due to thermal excitations.
Similarly, the observation of an infrared symmetric phase Alexandru:2019gdm around $T_{\mathrm{IR}}\sim 230$ MeV in $N_{f}=2+1$ lattice QCD Meng:2023nxf , which disappears by $T\gtrsim 300$ MeV, supports a correlated, quasi-hadronic medium in the window $T_{c}\lesssim T\lesssim 2T_{c}$. In this regime, hadrons become progressively delocalized, forming a dense fluid where both chromoelectric and chromomagnetic interactions remain active. The usual chiral multiplets are already restored in the nonsinglet sector, but chiral-spin symmetry is not yet manifest.
The gradual loss of binding across both regimes is consistent with the observed reduction in $U(1)_{A}$-breaking effects in connected channels (as measured by $\kappa_{PS}$ and $\kappa_{TX}$), while topological fluctuations—including center vortices tHooft:1977nqb and instanton-like objects—remain active until higher temperatures.
@@
These emergent symmetry patterns are complemented by topological studies. Center-vortex analyses in $N_{f}=2+1$ QCD observe a pronounced change in vortex percolation at approximately $\gtrsim 2T_{c}$ Mickley:2024vkm . While vortex density drops substantially, a sparse network survives to higher temperatures, reinforcing that topological fluctuations — intimately connected to the $U(1)_{A}$ anomaly — remain non-negligible well into the deconfined regime. This provides a plausible mechanism for continued $U(1)_{A}$ breaking where nonsinglet chiral symmetry is largely restored. The smooth crossover implies no sharp “complete deconfinement” point; instead, the system evolves gradually from hadronic to quark–gluon plasma, with residual correlations possible even at $T\gg 2T_{c}$.
@@
Taken together, these independent lines of evidence support the two-stage effective restoration scenario described in section 6.1. Stage 1 ($T\sim T_{c}^{ns}$) is characterized by effective restoration of nonsinglet symmetries, as demonstrated by the vanishing of $\kappa_{PS}$, $\kappa_{VA}$, and $\kappa_{TX}$ in this work. Stage 2 ($T\sim T_{c}^{s}$) involves full effective restoration of singlet channels and suppression of topological fluctuations ($\chi_{t}\to 0$). Consistent comparison between lattice studies requires continuum extrapolation of each observable, as demonstrated here. Since the chiral transition at physical quark masses is a smooth crossover, any symmetry-breaking observable decreases gradually with temperature. The Stage 2 scenario remains to be conclusively established through future continuum-extrapolated measurements of tensor-singlet ratios and topological susceptibility.
@@ 6.4.1 Computational improvements and tensor singlet probes
  • Flavor-singlet tensor channels: Using stochastic all-to-all propagators with color–Dirac dilution, we will compute the singlet tensor ($\omega_{T}$) and axial-tensor ($h_{1T}$) susceptibilities. The tensor singlet ratio
    $\kappa_{\omega_{T},h_{1T}}\equiv\frac{\chi_{\omega_{T}}-\chi_{h_{1T}}}{\chi_{\omega_{T}}+\chi_{h_{1T}}},$
    where $\omega_{T,k}=\bar{q}\gamma_{4}\gamma_{k}q$ and $h_{1T,k}=\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$, provides a theoretically clean (power-divergence free) probe of $U(1)_{A}$ restoration in the singlet sector. This will enable a direct test of the two-stage $U(1)_{A}$ restoration scenario at $T_{c}^{s}\gg T_{c}^{ns}$.
  • Mixed singlet–nonsinglet channels: The mixed ratios $\kappa(\chi_{\omega_{T}},\chi_{b_{1}})$ and $\kappa(\chi_{h_{1T}},\chi_{\rho_{T}})$ probe $SU(2)_{L}\times SU(2)_{R}$ restoration between singlet and nonsinglet tensor operators, offering additional insight into the chiral symmetry realization pattern across singlet and nonsinglet flavor sectors.